OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 + 3 √3i

OpenStudy (kc_kennylau):

One trick for you

OpenStudy (kc_kennylau):

\[r=\sqrt{a^2+b^2}\]

OpenStudy (kc_kennylau):

\[\theta=\tan^{-1}\left(\frac ba\right)\]

OpenStudy (anonymous):

Hmm how do I apply this?

OpenStudy (kc_kennylau):

Talking about a+bi ahaha

OpenStudy (kc_kennylau):

Never mind, you may have another method

OpenStudy (anonymous):

I got -60 for the tan-1 lol not sure what to do

OpenStudy (kc_kennylau):

You may wish to call it arctangent instead of tan-1, and the -60 should be in degrees

OpenStudy (kc_kennylau):

Make it positive

OpenStudy (anonymous):

Ah I'm confused now, sorry

OpenStudy (kc_kennylau):

Okay find r first and sorry for confusing you

OpenStudy (anonymous):

okay so r = √a^2+b^2 so √-3+3√3i right?

OpenStudy (kc_kennylau):

Well (1) you forgot to square and (2) the b doesn't include the i

OpenStudy (anonymous):

Ahh okay thanks! Let me try it now

OpenStudy (anonymous):

3.5

OpenStudy (kc_kennylau):

?

OpenStudy (anonymous):

I put it in my calculator and that's what I got lol

OpenStudy (kc_kennylau):

Put in your calculator \(\sqrt{3^2+(3\sqrt3)^2}\)

OpenStudy (anonymous):

Now its 4.2

OpenStudy (kc_kennylau):

Actually

OpenStudy (kc_kennylau):

Don't use your calculator

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So now what do I do?

OpenStudy (kc_kennylau):

Well calculate \(\sqrt{3^2+(3\sqrt3)^2}\)

OpenStudy (anonymous):

With my calculator? lol

OpenStudy (kc_kennylau):

lol without

OpenStudy (anonymous):

√12

OpenStudy (kc_kennylau):

Well it's \((3\sqrt3)^2\) not \((\sqrt3)^2\) inside

OpenStudy (anonymous):

Okay my calculator said something else but would it be 6? Because it would be √36?

OpenStudy (kc_kennylau):

\[\sqrt{3^2+(3\sqrt3)^2}=\sqrt{9+27}=\sqrt{36}=6\]

OpenStudy (kc_kennylau):

So back to the question

OpenStudy (anonymous):

Express it in trigonometric form

OpenStudy (kc_kennylau):

Yep

OpenStudy (kc_kennylau):

What's the trigonometric form?

OpenStudy (anonymous):

z=r(cosα+i⋅sinα)

OpenStudy (anonymous):

so the r is 6

OpenStudy (kc_kennylau):

Yes

OpenStudy (anonymous):

6(cosx+isinx)

OpenStudy (anonymous):

Now we need to find the cos x and sin x right

OpenStudy (kc_kennylau):

And I have to apologize for my accent that seems rude because I'm not in my full mood currently

OpenStudy (kc_kennylau):

The x is -60deg

OpenStudy (anonymous):

It's alright!! and okay how do I convert that to degrees again

OpenStudy (kc_kennylau):

-60deg is already degrees

OpenStudy (kc_kennylau):

The problem is how to make that positive

OpenStudy (anonymous):

Oh haha okay how do I do that

OpenStudy (kc_kennylau):

What's a full circle?

OpenStudy (anonymous):

360

OpenStudy (kc_kennylau):

Add it to that

OpenStudy (anonymous):

360+-60 = 300

OpenStudy (kc_kennylau):

Exactly

OpenStudy (kc_kennylau):

Back to the ques

OpenStudy (anonymous):

Trig form, okay hmm cos(300)?

OpenStudy (kc_kennylau):

Well

OpenStudy (kc_kennylau):

6(cos300+isin300)

OpenStudy (anonymous):

Ahh gotcha so now I have to relate it to the unit circle or something

OpenStudy (kc_kennylau):

Yep

OpenStudy (kc_kennylau):

The trigonometric enables you to plot the point on the Argand's diagram or however it's called coz I can't forking spell

OpenStudy (anonymous):

Okay so the point on the unit circle is 5pi/3 or (1/2, -√3/2)

OpenStudy (kc_kennylau):